USE ONE SAMPLE Z TEST


The average baptismal cost includes 50 guest. A random sample of 32 baptismal during the past year in the National Capital Region had a mean of 53 guest and a standard deviation of 10. Is there sufficient evidence at the 0.05 level of significance that the average number of guest differs from the national average?


Given:


STEP 1: STATE THE HYPOTHESIS AND IDENTIFY THE CLAIM.


STEP 2: THE LEVEL OF SIGNIFICANCE 


STEP 3: THE Z CRITICAL VALUE


STEP 4: COMPUTE THE ONE SAMPLE Z TEST VALUE

 

STEP 5: DECISION RULE


STEP 6: CONCLUSION


1
Expert's answer
2021-05-31T16:17:43-0400

Given:

"\bar{X}=53 \\

\sigma=10\\

n=32"

STEP 1: STATE THE HYPOTHESIS AND IDENTIFY THE CLAIM.

"H_0: \mu=50 \\

H_1: \mu ≠50"

STEP 2: THE LEVEL OF SIGNIFICANCE

α=0.05

STEP 3: THE Z CRITICAL VALUE

"Z=\frac{\bar{X}- \mu}{\sigma / \sqrt{n}}"

STEP 4: COMPUTE THE ONE SAMPLE Z TEST VALUE

"Z = \frac{53-50}{10/ \sqrt{32}}=\frac{3}{1.767}=1.697"

STEP 5: DECISION RULE

Two-tailed test.

Reject H0 if z≤-1.95 or z≥1.95.

STEP 6: CONCLUSION

Since Z=1.697 < 1.95

Thus, fail to reject the H0.

There is no sufficient evidence at the 0.05 level of significance that the average number of guest differs from the national average.